# Online, Target-Free LiDAR–Camera Extrinsic Calibration via Cross-Modal Mask Matching

### v1

@@ Abstract
LiDAR-camera extrinsic calibration (LCEC) is crucial for data fusion in intelligent vehicles. Offline, target-based approaches have long been the preferred choice in this field. However, they often demonstrate poor adaptability to real-world environments. This is largely because extrinsic parameters may change significantly due to moderate shocks or during extended operations in environments with vibrations. In contrast, online, target-free approaches provide greater adaptability yet typically lack robustness, primarily due to the challenges in cross-modal feature matching. Therefore, in this article, we unleash the full potential of large vision models (LVMs), which are emerging as a significant trend in the fields of computer vision and robotics, especially for embodied artificial intelligence, to achieve robust and accurate online, target-free LCEC across a variety of challenging scenarios. Our main contributions are threefold: we introduce a novel framework known as MIAS-LCEC, provide an open-source versatile calibration toolbox with an interactive visualization interface, and publish three real-world datasets captured from various indoor and outdoor environments. The cornerstone of our framework and toolbox is the cross-modal mask matching (C3M) algorithm, developed based on a state-of-the-art (SoTA) LVM and capable of generating sufficient and reliable matches. Extensive experiments conducted on these real-world datasets demonstrate the robustness of our approach and its superior performance compared to SoTA methods, particularly for the solid-state LiDARs with super-wide fields of view.
@@
LiDAR-camera extrinsic calibration, intelligent vehicle, large vision model, embodied artificial intelligence.figure Visualization of the experimental result achieved using our proposed online, targeless LCEC algorithm.
@@ I-A Background
LiDARs provide accurate spatial geometric information, while cameras capture rich textural details [1, 2, 3]. Fusing data from both sensors enables intelligent vehicles to achieve more comprehensive 3D environmental perception [4, 5, 6]. LiDAR-camera extrinsic calibration (LCEC) forms the foundation for this data fusion process [7, 8, 9], as illustrated in Fig. Online, Target-Free LiDAR–Camera Extrinsic Calibration via Cross-Modal Mask Matching. It basically estimates an extrinsic matrix , defined as follows [10]:
@@ I-B Existing Challenges and Motivation
On the other hand, point feature-based LCEC approaches require distinctive 2D image pixels and 3D LiDAR points, characterized by significant changes in intensity or depth across all dimensions. However, this dependency may lead to a scarcity of viable correspondences, especially in low-texture environments [18]. Additionally, these methods often overlook the alignment of the field of view (FoV) between LiDAR and camera, resulting in an excess of irrelevant points within the LiDAR point clouds, adversely impacting the LCEC process’s stability [17].
@@ I-C Novel Contributions
Therefore, in this article, we move one step forward in the field of online, target-free LCEC by unleashing the potential of MobileSAM [27], a SoTA LVM for image segmentation. First, we develop an online, target-free LCEC approach, referred to as MIAS-LCEC, which employs a novel coarse-to-fine strategy to accurately estimate LiDAR-camera extrinsic parameters. Specifically, a virtual camera is introduced with iterative pose updates to project the given LiDAR point cloud, generating a LiDAR intensity projection (LIP) image, which looks as if taken from the perspective of the actual camera. This addresses the oversight of FoV alignment in the prior study [17] and helps achieve more effective and robust 3D-2D correspondence matching. Subsequently, both the LIP and RGB images undergo segmentation using MobileSAM. These segmentation results are then processed using a novel cross-modal mask matching (C3M) algorithm, capable of generating sparse yet reliable matches, which are propagated to target masks for dense matching. Finally, the obtained correspondences serve as inputs for a perspective-n-points (PnP) solver to derive the extrinsic matrix. Additionally, we launch a powerful toolbox with an interactive visualization interface. This toolbox also incorporates functionalities for both offline, target-based calibration, and manual calibration, thereby further improving its utility. We collect three real-world datasets (from a variety of indoor and outdoor environments under various scenarios as well as different weather and illumination conditions) using a CMOS camera and different types of solid-state LiDARs to comprehensively evaluate the performance of LCEC algorithms. Through extensive experiments conducted on these datasets, our proposed MIAS-LCEC demonstrates superior robustness and accuracy compared to other SoTA online, target-free approaches. Moreover, it achieves a similar performance to that of an offline, target-based algorithm. Our source code, toolbox, and datasets will be publicly available upon publication at mias.group/MIAS-LCEC.
@@
A versatile LCEC toolbox with an interactive visualization interface and capable of conducting online, target-free calibration, offline, target-based calibration, and manual calibration.
@@ II-B Target-Free Approaches
Afterwards, traditional line/edge feature-based automatic LCEC approaches emerged. In studies such as [12, 15], LiDAR point intensities are first projected into the camera perspective, thereby generating a virtual image, namely an LIP image. Edges are then extracted from both the LIP and RGB images. By matching these cross-modal edges, the relative pose between the two sensors can be determined. Similarly, research by [37, 38] optimizes extrinsic calibration by maximizing the mutual information between LIP and RGB images. To address occlusion issues in [37, 38], the study of HKU-Mars [14] employs a voxelization method to detect and extract 3D lines from the point cloud, achieving high accuracy in scenarios with rich 3D line features. While effective in specific scenarios with abundant features, these traditional methods heavily rely on well-distributed line/edge features, which can compromise calibration robustness. Moreover, the use of low-level image processing algorithms, such as Gaussian blur and the Canny operator, can introduce errors in edge detection, potentially fragmenting global lines and thus reducing overall calibration accuracy.
Advances in deep learning techniques have driven significant exploration into enhancing traditional line/edge feature-based algorithms. In [11], edge dection with Transformer (EDTER) [39], a deep neural network, is utilized to improve the accuracy of 2D edge detection in RGB image. Additionally, a supervoxel-based 3D line detection method was designed to detect global line features from 3D point clouds, improving the line-based method proposed in [14]. Despite its impressive performance, this approach still heavily relies on specific scenarios with abundant properly-distributed lines, ultimately limiting its applicability.
@@ III-B Cross-Modal Mask Matching
The corner points along the contours of masks detected within the LIP image and the RGB image using MobileSAM are represented by two sets: and , respectively. An instance (bounding box), utilized to precisely fit around each mask, is centrally positioned at and has a dimension of pixels. To determine optimum instance matches, we construct a cost matrix , where the element at , namely:
denotes the matching cost between the -th instance from the LIP image and the -th instance from the RGB image. is initially set as during the sparse matching phase and subsequently updated using the above-mentioned affine transformation prior to the dense matching phase, as illustrated in Fig. 2, so as to minimize the discrepancies arising from the differing perspectives between LiDAR and camera. Matches with the lowest costs in both horizontal and vertical directions are determined as the optimum coarse instance matches.
@@
is the angle between the vectors originating from the mask centers and pointing to their respective matched corner points. can then be yielded as follows:
which represents the ratio between the areas of the bounding boxes associated with the RGB image and the LIP image. Finally, according to (8), can be yielded as follows:
@@
where and represent a affine transformation from to . Combining (14) and (13) results in the following expression:
@@ IV-A Experimental Setup
We compare our method with two SoTA target-free LCEC approaches: HKU-Mars [14] and DVL [17], on two real-world datasets, MIAS-LCEC-TL70 and MIAS-LCEC-TL360. Additionally, to validate the effectiveness of our method in scenarios where targets are present, we also conduct comparisons with a classical offline target-based LCEC approach introduced in [28] on the MIAS-LCEC-CB70 dataset.
Our algorithm was implemented on an Intel i7-12800K CPU (utilizing seven threads) and an NVIDIA RTX3080Ti GPU. The entire process, including data preprocessing, C3M, and extrinsic parameters optimization, requires approximately 15 seconds for one iteration and 70 seconds for three iterations.
@@
Residential Community | HKU-Mars [14] | 1.625 | 1.125 | 1.130 | 2.662 | 7.036 | 0.749 | 0.829 | 0.458 | 0.397 | 1.208 | 4.405 | 0.319 |
DVL [17] | 0.267 | 0.084 | 0.134 | 0.336 | 0.626 | 0.111 | 0.033 | 0.043 | 0.018 | 0.063 | 0.141 | 0.018 | |
MIAS-LCEC (Ours) | 0.199 | 0.124 | 0.103 | 0.305 | 0.711 | 0.012 | 0.032 | 0.017 | 0.026 | 0.050 | 0.176 | 0.018 | |
Urban Freeway | HKU-Mars [14] | 2.622 | 1.608 | 0.983 | 2.444 | 3.908 | 0.431 | 1.482 | 1.125 | 0.281 | 1.954 | 9.454 | 0.141 |
DVL [17] | 0.463 | 0.141 | 0.068 | 0.504 | 0.598 | 0.322 | 0.103 | 0.056 | 0.016 | 0.124 | 0.268 | 0.067 | |
MIAS-LCEC (Ours) | 0.458 | 0.110 | 0.095 | 0.491 | 0.697 | 0.312 | 0.070 | 0.052 | 0.010 | 0.111 | 0.144 | 0.045 | |
Building | HKU-Mars [14] | 1.515 | 1.447 | 0.596 | 2.169 | 3.727 | 0.923 | 0.533 | 0.262 | 0.231 | 0.708 | 2.596 | 0.065 |
DVL [17] | 0.959 | 0.081 | 0.079 | 0.968 | 1.174 | 0.866 | 0.058 | 0.047 | 0.037 | 0.090 | 0.148 | 0.056 | |
MIAS-LCEC (Ours) | 0.902 | 0.085 | 0.054 | 0.910 | 1.052 | 0.826 | 0.049 | 0.036 | 0.023 | 0.072 | 0.138 | 0.029 | |
Challenging Weather | HKU-Mars [14] | 1.135 | 1.153 | 1.783 | 2.579 | 11.232 | 0.444 | 0.772 | 0.456 | 0.471 | 1.086 | 4.930 | 0.303 |
DVL [17] | 0.162 | 0.132 | 0.162 | 0.309 | 0.517 | 0.086 | 0.029 | 0.036 | 0.020 | 0.055 | 0.113 | 0.021 | |
MIAS-LCEC (Ours) | 0.166 | 0.132 | 0.140 | 0.296 | 0.486 | 0.160 | 0.028 | 0.027 | 0.020 | 0.047 | 0.124 | 0.021 | |
Indoor | HKU-Mars [14] | 1.068 | 1.393 | 1.507 | 2.458 | 5.370 | 0.442 | 0.167 | 0.075 | 0.123 | 0.248 | 0.890 | 0.036 |
DVL [17] | 0.258 | 0.148 | 0.089 | 0.330 | 0.776 | 0.113 | 0.020 | 0.016 | 0.010 | 0.031 | 0.075 | 0.011 | |
MIAS-LCEC (Ours) | 0.193 | 0.158 | 0.170 | 0.343 | 0.730 | 0.158 | 0.013 | 0.011 | 0.012 | 0.024 | 0.042 | 0.011 | |
Challenging Illumination | HKU-Mars [14] | 7.755 | 8.719 | 7.200 | 14.909 | 38.634 | 0.366 | 1.611 | 1.578 | 1.678 | 3.386 | 10.523 | 0.035 |
DVL [17] | 0.551 | 1.550 | 0.291 | 1.727 | 8.579 | 0.089 | 0.329 | 0.144 | 0.101 | 0.378 | 1.974 | 0.019 | |
MIAS-LCEC (Ours) | 0.341 | 0.498 | 0.247 | 0.759 | 1.473 | 0.231 | 0.084 | 0.057 | 0.032 | 0.116 | 0.216 | 0.029 | |
All | HKU-Mars [14] | 2.218 | 2.156 | 1.967 | 3.954 | 38.634 | 0.366 | 0.807 | 0.555 | 0.476 | 1.262 | 10.523 | 0.035 |
DVL [17] | 0.378 | 0.283 | 0.136 | 0.582 | 8.579 | 0.086 | 0.074 | 0.050 | 0.029 | 0.101 | 1.974 | 0.011 | |
MIAS-LCEC (Ours) | 0.315 | 0.170 | 0.134 | 0.453 | 1.473 | 0.012 | 0.039 | 0.028 | 0.021 | 0.061 | 0.216 | 0.011 |

### v2

# Online, Target-Free LiDAR-Camera Extrinsic Calibration via Cross-Modal Mask Matching

@@ Abstract
LiDAR-camera extrinsic calibration (LCEC) is crucial for data fusion in intelligent vehicles. Offline, target-based approaches have long been the preferred choice in this field. However, they often demonstrate poor adaptability to real-world environments. This is largely because extrinsic parameters may change significantly due to moderate shocks or during extended operations in environments with vibrations. In contrast, online, target-free approaches provide greater adaptability yet typically lack robustness, primarily due to the challenges in cross-modal feature matching. Therefore, in this article, we unleash the full potential of large vision models (LVMs), which are emerging as a significant trend in the fields of computer vision and robotics, especially for embodied artificial intelligence, to achieve robust and accurate online, target-free LCEC across a variety of challenging scenarios. Our main contributions are threefold: we introduce a novel framework known as MIAS-LCEC, provide an open-source versatile calibration toolbox with an interactive visualization interface, and publish three real-world datasets captured from various indoor and outdoor environments. The cornerstone of our framework and toolbox is the cross-modal mask matching (C3M) algorithm, developed based on a state-of-the-art (SoTA) LVM and capable of generating sufficient and reliable matches. Extensive experiments conducted on these real-world datasets demonstrate the robustness of our approach and its superior performance compared to SoTA methods, particularly for the solid-state LiDARs with super-wide fields of view. Our toolbox and datasets are publicly available at mias.group/MIAS-LCEC.
@@
LiDAR-camera extrinsic calibration, intelligent vehicle, large vision model, embodied artificial intelligence.![[Uncaptioned image]](x1.png)
figureVisualization of the experimental results achieved using our proposed online, target-free LCEC algorithm.
@@ I-A Background
LiDARs provide accurate spatial geometric information, while cameras capture rich textural details [1, 2, 3, 4]. Fusing data from both sensors enables intelligent vehicles to achieve more comprehensive 3D environmental perception [5, 6, 7, 8, 9, 10]. LiDAR-camera extrinsic calibration (LCEC) forms the foundation for this data fusion process [11, 12, 13, 14], as illustrated in Fig. Online, Target-Free LiDAR-Camera Extrinsic Calibration via Cross-Modal Mask Matching. It basically estimates an extrinsic matrix , defined as follows [15]:
@@ I-B Existing Challenges and Motivation
On the other hand, point feature-based LCEC approaches require distinctive 2D image pixels and 3D LiDAR points, characterized by significant changes in intensity or depth across all dimensions. However, this dependency may lead to a scarcity of viable correspondences, especially in low-texture environments [23]. Additionally, these methods often overlook the alignment of the field of view (FoV) between LiDAR and camera, resulting in excess of irrelevant points within the LiDAR point clouds, adversely impacting the LCEC process’s stability [22].
@@ I-C Novel Contributions
Therefore, in this article, we move one step forward in the field of online, target-free LCEC by unleashing the potential of MobileSAM [32], a SoTA LVM for image segmentation. First, we develop an online, target-free LCEC approach, referred to as MIAS-LCEC, which employs a novel coarse-to-fine strategy to accurately estimate LiDAR-camera extrinsic parameters. To minimize the modality discrepancy, we formulate the 3D-2D feature matching as a 2D-2D feature matching problem by introducing a virtual camera (whose pose is iteratively updated) to project the given LiDAR point cloud, thereby generating a LiDAR intensity projection (LIP) image, which appears as if it were taken from the perspective of the actual camera. This addresses the oversight of FoV alignment in the prior study [22] and helps achieve more effective and robust 3D-2D correspondence matching. Subsequently, both the LIP and RGB images undergo segmentation using MobileSAM. These segmentation results are then processed using a novel cross-modal mask matching (C3M) algorithm, capable of generating sparse yet reliable matches, which are propagated to target masks for dense matching. Finally, the obtained correspondences serve as inputs for a Perspective-n-Point (PnP) solver to derive the extrinsic matrix. Additionally, we launch a powerful toolbox with an interactive visualization interface. This toolbox also incorporates the manual calibration functionality, thereby further improving its utility. We collect three real-world datasets (from a variety of indoor and outdoor environments under various scenarios as well as different weather and illumination conditions) using a CMOS camera and different types of solid-state LiDARs to comprehensively evaluate the performance of LCEC algorithms. Through extensive experiments conducted on these datasets, our proposed MIAS-LCEC demonstrates superior robustness and accuracy compared to other SoTA online, target-free approaches. Moreover, it achieves a similar performance to that of an offline, target-based algorithm. Our toolbox and datasets are publicly available at mias.group/MIAS-LCEC.
@@
A versatile LCEC toolbox with an interactive visualization interface and capable of conducting online, target-free calibration and manual calibration.
@@ II-B Target-Free Approaches
Afterwards, traditional line/edge feature-based automatic LCEC approaches emerged. In studies such as [17, 20], LiDAR point intensities are first projected into the camera perspective, thereby generating a virtual image, namely an LIP image. Edges are then extracted from both the LIP and RGB images. By matching these cross-modal edges, the relative pose between the two sensors can be determined. Similarly, research by [42, 43] optimizes extrinsic calibration by maximizing the mutual information (MI) between LIP and RGB images. To address occlusion issues in [42, 43], the study of HKU-Mars [19] employs a voxelization method to detect and extract 3D lines from the point cloud, achieving high accuracy in scenarios with rich 3D line features. While effective in specific scenarios with abundant features, these traditional methods heavily rely on well-distributed line/edge features, which can compromise calibration robustness. Moreover, the use of low-level image processing algorithms, such as Gaussian blur and the Canny operator, can introduce errors in edge detection, potentially fragmenting global lines and thus reducing overall calibration accuracy.
Advances in deep learning techniques have driven significant exploration into enhancing traditional line/edge feature-based algorithms. In [16], edge detection with Transformer (EDTER) [44], a deep neural network, is utilized to improve the accuracy of 2D edge detection in RGB images. Additionally, a supervoxel-based 3D line detection method was designed to detect global line features from 3D point clouds, improving the line-based method proposed in [19]. Despite its impressive performance, this approach still heavily relies on specific scenarios with abundant properly-distributed lines, ultimately limiting its applicability.
@@ III-B Cross-Modal Mask Matching
As shown in Fig. 2, our developed C3M strategy significantly minimizes potential matching errors produced by the LVM, primarily due to the strict match selection criterion. The corner points along the contours of masks detected within the LIP image and the RGB image using MobileSAM are represented by two sets: and , respectively. An instance (bounding box), utilized to precisely fit around each mask, is centrally positioned at and has a dimension of pixels. To determine optimum instance matches, we construct a cost matrix , where the element at , namely:
denotes the matching cost between the -th instance from the LIP image and the -th instance from the RGB image. is initially set as during the sparse matching phase and subsequently updated using the above-mentioned affine transformation prior to the dense matching phase, as illustrated in Fig. 3, so as to minimize the discrepancies arising from the differing perspectives between LiDAR and camera. A strict criterion is applied to achieve sparse yet reliable matching. Matches with the lowest costs in both horizontal and vertical directions are determined as the optimum coarse instance matches.
@@
is the angle between the vectors originating from the mask centers and pointing to their respective matched corner points. can then be expressed as follows:
which represents the ratio between the areas of the bounding boxes associated with the RGB image and the LIP image. Finally, according to (8), can be obtained as follows:
@@
where and represent an affine transformation from to . Combining (14) and (13) results in the following expression:
@@ IV-A Experimental Setup
We compare our method with four SoTA target-free LCEC approaches: CRLF [25], UMich [43], HKU-Mars [19] and DVL [22], on two real-world datasets, MIAS-LCEC-TF70 and MIAS-LCEC-TF360. Additionally, to validate the effectiveness of our method in scenarios where targets are present, we also conduct comparisons with a classical offline target-based LCEC approach introduced in [33] on the MIAS-LCEC-CB70 dataset.
Our algorithm was implemented on an Intel i7-14700K CPU and an NVIDIA RTX4070Ti Super GPU. The entire process, including data preprocessing, C3M, and extrinsic parameters optimization, takes approximately 15 to 70 seconds.
@@
Residential Community | CRLF [25] | 0.177 | 1.581 | 0.876 | 1.594 | 1.754 | 1.537 | 0.162 | 0.429 | 0.060 | 0.464 | 3.636 | 0.136 |
UMich [43] | 0.690 | 1.502 | 4.066 | 4.829 | 21.706 | 0.214 | 0.306 | 0.151 | 0.058 | 0.387 | 2.394 | 0.072 | |
HKU-Mars [19] | 1.704 | 1.128 | 1.113 | 2.695 | 7.036 | 0.749 | 0.829 | 0.458 | 0.397 | 1.208 | 4.405 | 0.319 | |
DVL [22] | 0.107 | 0.089 | 0.091 | 0.193 | 0.383 | 0.042 | 0.034 | 0.043 | 0.018 | 0.063 | 0.141 | 0.018 | |
MIAS-LCEC (Ours) | 0.049 | 0.138 | 0.094 | 0.190 | 0.711 | 0.012 | 0.032 | 0.016 | 0.026 | 0.050 | 0.176 | 0.018 | |
Urban Freeway | CRLF [25] | 0.434 | 1.584 | 0.857 | 1.582 | 1.585 | 1.581 | 0.043 | 0.130 | 0.030 | 0.140 | 0.140 | 0.140 |
UMich [43] | 0.471 | 1.331 | 1.464 | 2.267 | 6.531 | 0.506 | 0.051 | 0.128 | 0.063 | 0.166 | 0.241 | 0.103 | |
HKU-Mars [19] | 2.518 | 1.610 | 0.955 | 2.399 | 3.744 | 0.337 | 1.482 | 1.125 | 0.284 | 1.956 | 9.453 | 0.148 | |
DVL [22] | 0.212 | 0.140 | 0.045 | 0.298 | 0.420 | 0.090 | 0.103 | 0.056 | 0.017 | 0.124 | 0.268 | 0.064 | |
MIAS-LCEC (Ours) | 0.216 | 0.111 | 0.083 | 0.291 | 0.514 | 0.062 | 0.070 | 0.052 | 0.010 | 0.111 | 0.145 | 0.047 | |
Building | CRLF [25] | 0.231 | 1.409 | 0.911 | 1.499 | 1.706 | 1.465 | 17.078 | 8.692 | 6.168 | 20.165 | 140.316 | 0.140 |
UMich [43] | 0.927 | 1.824 | 15.097 | 11.914 | 22.778 | 0.452 | 0.544 | 0.182 | 0.407 | 0.781 | 1.497 | 0.042 | |
HKU-Mars [19] | 0.800 | 1.391 | 0.589 | 1.814 | 3.618 | 0.118 | 0.532 | 0.262 | 0.227 | 0.706 | 2.595 | 0.059 | |
DVL [22] | 0.138 | 0.103 | 0.049 | 0.200 | 0.357 | 0.059 | 0.058 | 0.047 | 0.031 | 0.087 | 0.146 | 0.052 | |
MIAS-LCEC (Ours) | 0.080 | 0.152 | 0.074 | 0.198 | 0.390 | 0.051 | 0.049 | 0.036 | 0.020 | 0.072 | 0.137 | 0.024 | |
Challenging Weather | CRLF [25] | 0.183 | 1.647 | 0.917 | 1.646 | 1.803 | 1.552 | 1.614 | 0.867 | 0.617 | 2.055 | 20.307 | 0.137 |
UMich [43] | 0.383 | 1.635 | 3.720 | 1.851 | 5.335 | 0.294 | 0.212 | 0.109 | 0.103 | 0.310 | 2.134 | 0.043 | |
HKU-Mars [19] | 1.110 | 1.128 | 1.840 | 2.578 | 11.647 | 0.371 | 0.772 | 0.456 | 0.469 | 1.086 | 4.934 | 0.302 | |
DVL [22] | 0.095 | 0.100 | 0.069 | 0.181 | 0.311 | 0.037 | 0.029 | 0.036 | 0.015 | 0.052 | 0.113 | 0.017 | |
MIAS-LCEC (Ours) | 0.100 | 0.112 | 0.059 | 0.177 | 0.412 | 0.030 | 0.027 | 0.027 | 0.018 | 0.046 | 0.127 | 0.021 | |
Indoor | CRLF [25] | 0.164 | 1.721 | 1.010 | 1.886 | 2.201 | 1.551 | 27.895 | 6.880 | 6.368 | 30.046 | 90.926 | 0.140 |
UMich [43] | 0.135 | 1.531 | 0.971 | 2.029 | 4.596 | 0.448 | 0.062 | 0.062 | 0.040 | 0.109 | 0.176 | 0.020 | |
HKU-Mars [19] | 1.065 | 1.414 | 1.572 | 2.527 | 5.458 | 0.779 | 0.165 | 0.075 | 0.122 | 0.246 | 0.893 | 0.036 | |
DVL [22] | 0.261 | 0.194 | 0.129 | 0.391 | 0.879 | 0.193 | 0.022 | 0.016 | 0.007 | 0.030 | 0.078 | 0.013 | |
MIAS-LCEC (Ours) | 0.144 | 0.234 | 0.175 | 0.363 | 0.724 | 0.225 | 0.016 | 0.011 | 0.011 | 0.024 | 0.045 | 0.012 | |
Challenging Illumination | CRLF [25] | 0.088 | 1.778 | 1.156 | 1.876 | 2.141 | 1.613 | 18.080 | 2.954 | 3.886 | 19.047 | 34.444 | 0.132 |
UMich [43] | 0.361 | 4.841 | 2.924 | 5.012 | 12.087 | 0.314 | 0.108 | 0.223 | 0.148 | 0.330 | 0.679 | 0.043 | |
HKU-Mars [19] | 7.762 | 8.774 | 7.235 | 14.996 | 38.750 | 0.338 | 1.609 | 1.578 | 1.679 | 3.386 | 10.520 | 0.034 | |
DVL [22] | 0.546 | 1.573 | 0.256 | 1.747 | 8.466 | 0.207 | 0.328 | 0.144 | 0.099 | 0.377 | 1.970 | 0.022 | |
MIAS-LCEC (Ours) | 0.347 | 0.524 | 0.180 | 0.749 | 1.554 | 0.149 | 0.087 | 0.057 | 0.030 | 0.118 | 0.220 | 0.033 | |
All | CRLF [25] | 0.197 | 1.625 | 0.946 | 1.683 | 2.201 | 1.465 | 10.051 | 3.036 | 2.589 | 11.133 | 140.316 | 0.132 |
UMich [43] | 0.485 | 1.945 | 4.272 | 4.265 | 22.778 | 0.214 | 0.217 | 0.134 | 0.115 | 0.333 | 2.394 | 0.020 | |
HKU-Mars [19] | 2.140 | 2.156 | 1.988 | 3.941 | 38.750 | 0.118 | 0.806 | 0.555 | 0.475 | 1.261 | 10.520 | 0.034 | |
DVL [22] | 0.201 | 0.292 | 0.104 | 0.423 | 8.466 | 0.037 | 0.075 | 0.050 | 0.026 | 0.100 | 1.970 | 0.013 | |
MIAS-LCEC (Ours) | 0.133 | 0.196 | 0.110 | 0.298 | 1.554 | 0.012 | 0.040 | 0.028 | 0.019 | 0.061 | 0.220 | 0.012 |